OpenStudy (anonymous):

what is -5x+2y=20 in slope intercept form

OpenStudy (anonymous):

It is in slope -intercept form only

OpenStudy (anonymous):

Any equation of the form y = mx + b is said to be in slope intercept form

OpenStudy (anonymous):

Just u need to rearrange the terms 2y = 20 + 5x y = 10 + 5/2x

OpenStudy (anonymous):

A science teacher has a supply of 50% sugar solution and a supply of 80% sugar solution. How much of each solution should the teacher mix together to get 150 mL of 60% sugar solution? a. 70 mL of the 50% solution and 35mL of the 80% solution. n. 35mL of the 50% solution and 70mL of the 80% solution. c. 70mL of the 50% solution and 70mL of the 80% solution. d. 35mL of the 50% solution and 35mL of the 80% solution.